OpenStudy (anonymous):

What is an amount of money multiplied by the interest rate and the amount of time that the money will be earning interest?

OpenStudy (anonymous):

Not entirely clear what your question is. Usually we call the starting amount of money the "Present Value" (PV for short). So if I have $100 today my present value would be $100. Often we want to see how much money we will have in the future if we can invest our money at a given interest rate. This value in known as the future value (FV). We also need to know what kind of interest rate we will be earning. So to figure out our future value for a given interest rate: FV=PF(1+I)^t. t represents how long (usually years) you are going to save your money. So example, lets say I have a $100 invested earning 8% interest and I will invest it for 5 years. We would have FV=100*(1+0.08)^5=$146.93

OpenStudy (anonymous):

to close things up...this process in general is known as compounding.